OpenStudy (anonymous):

Compute a value for t that satisfies the equation below. -2/3t - 3 =-5

OpenStudy (anonymous):

sir do you know the OS rules

OpenStudy (anonymous):

is this ur equation ? \(\Huge\frac{-2}{3t-3}=-5\)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-2/3t

OpenStudy (anonymous):

- 3

OpenStudy (anonymous):

= 5

OpenStudy (anonymous):

this one ? \(\Huge\frac{-2}{3t}-3=-5\)

OpenStudy (anonymous):

yeah ^-^ how do you do that exactly ?

OpenStudy (anonymous):

like how did you write that ?

OpenStudy (anonymous):

its \(\Huge\color{green}{LATEX} \) so , first start with adding +3 for both sides , like this \(\Huge\frac{-2}{3t}-3+3=-5+3\)

OpenStudy (anonymous):

\(\Huge\frac{-2}{3t} =-2\) got it untill now ?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

then do the \(\Huge\color{blue}{Cross-multiplication}\) like this \(\Huge -2=-2×3t\)

OpenStudy (anonymous):

so as u can see you would have this equation \(\Huge -6t=-2\) can u find t now ?

OpenStudy (anonymous):

yes thank you ^-^

OpenStudy (anonymous):

then devide by \( \color{red}{-6}\) \(\Huge\color{green}{t}=\frac{-2}{-6}=\frac{1}{3}\)

OpenStudy (anonymous):

so 3?

OpenStudy (anonymous):

no , check again , its : \(\Huge\color{red}{\frac{1}{3}}\)

OpenStudy (anonymous):

and ur wlc any time ^_^

OpenStudy (anonymous):

ohhhhh okay thanks again ^-^

OpenStudy (anonymous):

np ur wlc ^_^